# SpanTrain: Highly Efficient Cross-domain Model Distributed Training System under Heterogeneous GPUs and Networks in CEE Environment

### v1

@@ 4.4. Evaluating SpanTrain’s Adaptation
Figure 12. Evaluation results for SpanTrain’s Adapter component.
@@
Evaluation results for SpanTrain’s Adapter component.
As shown in Figure 12b, with Setting 1, we evaluate the SpanTrain’s adaption in the network fluctuation environment. As shown in Figure 12a, the ResNet101 and GPT-2 models are divided into 4 stages and spread over different 4 first-level network device groups. During our experiments, we add the noise to the network between 4 first-level network device groups, reducing their bandwidth by 40-60%, to emulate the network fluctuation. We evaluate four scenarios formed by two binary conditions: (1) Adapter activated or not and (2) network noise added or not. Specifically, including ① with the Adapter off and with network bandwidth degradation (FT), ② with the Adapter off and without network bandwidth degradation (FF), ③ with the Adapter on and with network bandwidth degradation (TT), and ④ with the Adapter on and without network bandwidth degradation (TF).
@@ 4.5. Scalability Evaluation
As illustrated in Figure 13b, the scalability of SpanTrain is evaluated by examining the model training throughput under varying numbers of devices. Based on the above configuration of the CEE, we ran this experiment by setting the server number of edge and end to be 1, 3, and 6, respectively, with the number of cloud servers being fixed, and using the network Setting 1 configuration described in §4.1. ResNet101 and GPT-2 models are used to be trained, and the batch size is set to be 128.
The experimental results show that SpanTrain’s average training throughput outperforms that of PyTorch-FSDP and GPipe by 2.79$\times$ and 1.64$\times$, respectively. That is because PyTorch FSDP requires weight synchronization across multiple nodes, where larger node counts exacerbate communication bottlenecks from stragglers. Meanwhile, GPipe’s 1F1B pipeline structure suffers from cascading inefficiencies: as model stages increase, (1) slow networks affect more stages, and (2) GPU bubble time grows proportionally with device count, significantly limiting throughput growth. In contrast, SpanTrain’s compact zero-bubble pipeline achieves superior scaling by: (1) decoupling weight updates to scheduling, and (2) bubble-padded execution that maximizes device resource utilization. Our approach reduces GPU idle time and demonstrates better performance scaling with increasing devices compared to existing methods.
@@ 4.6. System Overhead
Next, we evaluate SpanTrain’s overhead imposed by the Profiler, Planner, and Adapter components, respectively.
@@
Overall, these are shown to be that the time and storage overhead imposed by SpanTrain is lightweight and causes minimal impact in model training.
@@ 6. Conclusion
This paper proposes SpanTrain, a geo-distributed model training system, which is a communication-centric training framework to tackle the issue of slow and unstable inter-region networks in the cross-region training. Specifically, SpanTrain can provide a compact, zero-bubble pipeline parallelism, with automatically deriving optimal parallel strategies, and react to network fluctuations in real time. The experimental results show that SpanTrain can provide an efficient training performance in contrast to the SOTA methods and existing parallel training frameworks in the CEE environment.

### v2

@@ 4.4. Evaluating DeepCEE’s Adaptation
Figure 12. Evaluation results for DeepCEE’s Adapter component.
@@
Evaluation results for DeepCEE’s Adapter component.
As shown in Figure 12b, with Setting 1, we evaluate the DeepCEE’s adaption in the network fluctuation environment. As shown in Figure 12a, the ResNet101 and GPT-2 models are divided into 4 stages and spread over different 4 first-level network device groups. During our experiments, we add the noise to the network between 4 first-level network device groups, reducing their bandwidth by 40-60%, to emulate the network fluctuation. We evaluate four scenarios formed by two binary conditions: (1) Adapter activated or not and (2) network noise added or not. Specifically, including ① with the Adapter off and with network bandwidth degradation (FT), ② with the Adapter off and without network bandwidth degradation (FF), ③ with the Adapter on and with network bandwidth degradation (TT), and ④ with the Adapter on and without network bandwidth degradation (TF).
@@ 4.5. Scalability Evaluation
As illustrated in Figure 13b, the scalability of DeepCEE is evaluated by examining the model training throughput under varying numbers of devices. Based on the above configuration of the CEE, we ran this experiment by setting the server number of edge and end to be 1, 3, and 6, respectively, with the number of cloud servers being fixed, and using the network Setting 1 configuration described in §4.1. ResNet101 and GPT-2 models are used to be trained, and the batch size is set to be 128.
The experimental results show that DeepCEE’s average training throughput outperforms that of PyTorch-FSDP and GPipe by 2.79$\times$ and 1.64$\times$, respectively. That is because PyTorch FSDP requires weight synchronization across multiple nodes, where larger node counts exacerbate communication bottlenecks from stragglers. Meanwhile, GPipe’s 1F1B pipeline structure suffers from cascading inefficiencies: as model stages increase, (1) slow networks affect more stages, and (2) GPU bubble time grows proportionally with device count, significantly limiting throughput growth. In contrast, DeepCEE’s compact zero-bubble pipeline achieves superior scaling by: (1) decoupling weight updates to scheduling, and (2) bubble-padded execution that maximizes device resource utilization. Our approach reduces GPU idle time and demonstrates better performance scaling with increasing devices compared to existing methods.
@@ 4.6. System Overhead
Next, we evaluate DeepCEE’s overhead imposed by the Profiler, Planner, and Adapter components, respectively.
@@
Overall, these are shown to be that the time and storage overhead imposed by DeepCEE is lightweight and causes minimal impact in model training.
@@ 6. Conclusion
This paper proposes DeepCEE, a geo-distributed model training system, which is a communication-centric training framework to tackle the issue of slow and unstable inter-region networks in the cross-region training. Specifically, DeepCEE can provide a compact, zero-bubble pipeline parallelism, with automatically deriving optimal parallel strategies, and react to network fluctuations in real time. The experimental results show that DeepCEE can provide an efficient training performance in contrast to the SOTA methods and existing parallel training frameworks in the CEE environment.